Example 10-5 Consider the simple horizontal curve given in Example 10-1, with a tangent distance T= 229.27 ft and an arc length L 391.53 ft. If the station of the PI is established as 7 +47.64, find the stations of the PC and the PT
Answer #1

Given,

Tangent distance, T = 229.27 ft

Length of curve, L = 391.53 ft

Station of PI = 7 + 47.64

Station of PC = Station of PI - (L/2) = (7+47.64) - (1+95.77) = (5+51.87)

Station of PT = Station of PI + (L/2) = (7+47.64) + (1+95.77) = (9+43.41)